### v1

# Clickbait Detection via Large Language Models

Abstract
Clickbait, which aims to induce users with some surprising and even thrilling headlines for increasing click-through rates, permeates almost all online content publishers, such as news portals and social media. Recently, Large Language Models (LLMs) have emerged as a powerful instrument and achieved tremendous success in a serious of NLP downstream tasks. However, it is not yet known whether LLMs can be served as a high-quality clickbait detection system. In this paper, we analyze the performance of LLMs in the few-shot scenarios on a number of English and Chinese benchmark datasets. Experimental results show that LLMs cannot achieve the best results compared to the state-of-the-art deep and fine-tuning PLMs methods. Different from the human intuition, the experiments demonstrated that LLMs cannot make satisfied clickbait detection just by the headlines.
1 Introduction
With the rapid development of online applications, some content publishers try to utilize clickbait for generating profits Chen et al. (2015). Clickbait refers to deliberately enticed users to click with some curious and chilling headlines, which are always unrelated to the real content or even the advertising promotion Chakraborty et al. (2016). The popularity of clickbait will inevitably lead to the experience degradation or even the disgust of users, there is an urgent demand to develop effective automatic clickbait detection methods Liu et al. (2022).
In recent years, the research methods for clickbait detection evolved from feature engineering to neural networks and, more recently, into pre-trained language models. Feature engineering methods extracted features such as semantic and linguistic features for detection tasks Biyani et al. (2016); Wei and Wan (2017). The deep neural networks methods can learn more abstract and higher-level features by disentangling explanatory factors of variations behind news title and content for clickbait detection Yoon et al. (2019); Zheng et al. (2021). In recent years, Pre-trained Language Models (PLMs), such as BERT Devlin et al. (2018), also show superiority in clickbait detection task. However, the feature engineering-based methods and deep neural networks typically require large-scaled labeled data since the detection is regarded as a classification method. In the PLMs methods, the huge gap between pre-training and fine-tuning prevent detection tasks from fully utilizing pre-training knowledge.
More recently, Large Language Models (LLMs) have demonstrated the powerful ability in various NLP downstream tasks Brown et al. (2020); Chowdhery et al. (2022); Thoppilan et al. (2022), which can achieve awesome performance even in the few-shot and zero-shot scenarios. Nevertheless, it remains unclear how LLMs perform in clickbait detection tasks compared to current methods. To address this issue, in this paper, we conduct a systematic evaluation of the few-shot and zero-shot learning capability of LLMs. The experiments are validated on both English and Chinese open datasets, and we conducted an empirical comparison of the performance between ChatGPT with the GPT3.5 model (gpt-3.5-turbo) and other state-of-the-art methods such as prompt-tuning.
To the best of our knowledge, this is the first attempt to validate the performance of LLMs on clickbait detection, the works in this paper provide a preliminary evaluation including detection results and robustness. The key findings and insights are summarized as follows:
  • ChatGPT achieves unsatisfying results in the few-shot scenarios compared to the state-of-the-art clickbait detection methods. We found that fine-tuning PLMs and prompt-tuning can achieve better results just based on the titles, which is consistent with humans in the real-world.
  • ChatGPT is a monolithic model capable of supporting multiple languages, which makes it a comprehensive multilingual clickbait detection technique. After evaluating the performance of ChatGPT on the task of clickbait detection across two languages (English and Chinese), we observed that it achieved stable results on almost all evaluation metrics. This also confirms that LLMs can be adapted to other languages.
It is worth mentioning that the source code and all the results of this paper is available on https://github.com/zhuyiYZU/chatGPTforClickbait.
2 Related Work
2.1 Clickbait Detection
Clickbait detection is an emerging field which is attracting increasing attention in recent years. On most online services, such as e-commerce, social media, and news portals, more clicks means more profit and commercial revenue. Early clickbait detection methods mainly focus on extracting a variety of features for detection tasks, such as semantic Rony et al. (2017), linguistics Blom and Hansen (2015), and multi-modal features Chen et al. (2015). However, these methods require expert knowledge for feature selection, and the handcrafted features are limited in representing more abstract and higher-level information.
In recent years, deep neural network models facilitate crossing and combination among even more diverse and sophisticated features, which has shown fairly good performance in clickbait detection. The popular deep neural networks, such as Recurrent Neural Networks (RNN) Anand et al. (2017), Convolutional Neural Networks (CNN) Agrawal (2016), Attention Mechanism Mishra et al. (2020), and Graph Attention Networks Liu et al. (2022), have already been devoted to the clickbait detection tasks. Despite the success of deep learning methods, due to the requirements on large-scale labeled training datasets, lead to high costs in collecting eligible training data. Recently, some domain adaptation López-Sánchez et al. (2018) and data augmentation Yoon et al. (2019) methods have been proposed to address the issues, however, these methods may bring additional noise in detection tasks.
More recently, pre-trained language models (PLMs) such as BERT Devlin et al. (2018), RoBERTa Liu et al. (2019), and T5 Raffel et al. (2020) have emerged as a powerful instruments for language understanding and generation. Through the fine-tuning PLMs on the special downstream task, the rich knowledge distributed in PLMs can be stimulated to better serve downstream tasks including clickbait detection Indurthi et al. (2020); Yi et al. (2022). Despite the success of fine-tuning PLMs, some recent studies find one of its critical challenges is the significant gap of objective forms in pre-training and fine-tuning, which restricts taking full advantage of knowledge in PLMs.
2.2 Large Language Models
Represented by GPT-3 Brown et al. (2020), Large Language Models (LLMs) have achieved superior performance, especially in few-shot learning scenarios Brown et al. (2020); Chowdhery et al. (2022); Thoppilan et al. (2022). Different from the previous PLMs methods, LLMs has two distinct advantages. The first is the larger scale, LLMs have a much larger scale in terms of model parameters and training data. Secondly, without the fine-tuning PLMs, LLMs can prompt few-shot learning that requires no additional neural layer and shows excellent performance. However, there is no work on the capabilities of LLMs on clickbait detection tasks.
3 Methodology
Through the few-shot learning ability of LLMs, they can also achieve excellent performance for tasks with low resources. Considering the lack of large-scale training corpus for clickbait detection tasks, we will test the performance of LLMs in few-shot scenarios on clickbait detection.
LLMs typically use prompts (i.e. specific templates) to guide the model in predicting output or answers, without requiring specific training on the data. Utilizing this form of prompt, we conducted experiments with different prompts on OpenAI’s largest available model GPT3.5 (gpt-3.5-turbo) for ChatGPT.
Prompt. To stimulate the rich knowledge distributed in LLMs, we manually designed the prompts for validating the performance of clickbait detection. The details of the prompts are illustrated in Table 1. For the two prompts, {Clickbait Sentence} and {Not-Clickbait Sentence} refer to the input sentence, {Yes, it is a clickbait} and {No, it is not a clickbait} are the corresponding labels respectively. {Results} is the place the carries the outputs of LLMs.
Table 1: The hand-crafted prompts for clickbait detection.
Prompts
P1
I want you to detect whether
the input sentence is clickbait or not.
Input: {Clickbait Sentence}
Output: {Yes, it is a clickbait}
Input: {Not-Clickbait Sentence}
Output: {No, it is not a clickbait}
Input: {Clickbait or Not-Clickbait Sentence}
Output: {Results}
P2 Sentence:{Clickbait Sentence}
Question: Detect the above sentence is clickbait or not.
Answer: {Yes, it is a clickbait}
Sentence:{Not-Clickbait Sentence}
Question: Detect the above sentence is clickbait or not.
Answer: {No, it is not a clickbait}
Sentence:{Clickbait or Not-Clickbait Sentence}
Question: Detect the above sentence is clickbait or not.
Answer: {Results}
In the first prompt (P1), the {guide-input-output} pattern is employed to guide LLMs for detecting clickbaits. In the second prompt (P2), the {sentence-question-answer} pattern is utilized to detect clickbaits in the form of a question. It is worth mentioning that the specialized guide word, such as ”Output:” and ”Answer:”, are added to the prompts, then the clickbait detection is regarded as the close-style task for LLMs, which ensured to achievement of a unique output. When performing multilingual clickbait detection tasks, we translate these two prompts into Chinese used in specific tasks as shown in Table 2.
Table 2: The hand-crafted prompts for clickbait detection.
提示
P1
请帮我检测
以下输入是或不是点击诱饵。
输入: {点击诱饵句子}
输出:{是,这是一个点击诱饵}
输入: {非点击诱饵句子}
输出: {不,这不是一个点击诱饵}
输入: {点击诱饵或非点击诱饵句子}
输出: {结果}
P2 句子:{点击诱饵句子}
问题: 检测上面的句子是或不是点击诱饵。
回答: {是,这是一个点击诱饵}
句子:{非点击诱饵句子}
问题: 检测上面的句子是或不是点击诱饵。
回答: {不,这不是一个点击诱饵}
句子:{点击诱饵或非点击诱饵句子}
问题: 检测上面的句子是或不是点击诱饵。
回答: {结果}
Zero-shot. In the clickbait detection experiments in zero-shot scenarios, just one {guide-input-output} or {sentence-question-answer} pattern is directly input into the LLMs for the detection result.
Few-shot. In the few-shot scenario, some instances follow the {guide-input-output} or {sentence-question-answer} pattern are provided as the training data for the detection tasks. Notably, in the first prompt (P1), the guidance is not needed to be repeated, only the input-output is repeated for some instance sentences and corresponding labels are stacked. In the second prompt (P2), the sentence-question-answer is repeated for each sample as training data.
(a) English P1
(b) Chinese P1
(c) English P2
(d) Chinese P2
Figure 1: The example of ChatGPT detecting clickbait errors.
4 Experiment
4.1 Datasets and Templates
To evaluate the performance of our method for clickbait detection in both English and Chinese, we conduct experiments on seven datasets. The DL-Clickbait (DLC), Clickbait news detection (CND), SC-Clickbait (SCC) are three well-known public clickbait detection datasets in English. Sina, Tencent, Wechat, and Paper are four public clickbait detection datasets in Chinese. The statistical details of all the datasets are presented in Table 3.
Table 3: Statistics of the news datasets. ”#” and ”avg.#” denote ”the number of” and ”the average number of”.
Statistic DLC CND SCC Sina Tencent Wechat Paper
#total headline 2345 8324 32000 1912 2546 2800 30800
#train sets 1638 5801 22400 1338 1782 1960 21560
#test sets 707 2523 9600 574 764 840 9240
avg.# words per sentence 11 13 9 21 25 21 22
4.2 Baselines
We compare ChatGPT with the following deep neural network and fine-tuning PLMs methods.
TextCNN Agrawal (2016) TextCNN utilizes kernels with different sizes to find various features from the input text. These features are automatically detected and are used to train a neural network for the corresponding task. It understands the input text from different perspectives, and can be applied to various nature language processing tasks.
BiLSTM Kumar et al. (2018) A bidirectional LSTM with an attention mechanism to learn the extent to which a word contributes to the clickbait score of a social media post in a different way. It used a Siamese net to detect similarities between the source and target data. To add another layer of complexity to the model, it also use TextCNN to learn image embeddings from large amounts of data.
Gated_CNN Yang et al. (2019) The method can be stacked to represent large context sizes and extract hierarchical features over larger and larger contexts with more abstractive features.
FAST_TEXT Joulin et al. (2016) FAST_TEXT is a text-classification method, and the sequence of words is considered in it, which is a model learning distributed representations of words based on ordered words.
BERT Devlin et al. (2018) The BERT method uses words and sentences to distinguish the context of words in a sentence.
4.3 Implementation details and evaluation metrics
Few-shot settings
In this experiment, we randomly selected K(5,10,20) instances as the training set for ChatGPT, and part of the samples are selected as the test set. Considering that too few training samples may greatly affect the effectiveness of the baselines, we draw different training samples from different datasets, where the number of training samples is 200/300/600, 300/600/1200, and 500/1000/2000 for English datasets (DL-Clickbait, CND, and SC-Clickbait datasets) and 200/400/800, 400/800/1600, 400/800/1600 and 800/1600/3200 for Chinese datasets (Sina, Tencent, Wechat and Paper), corresponding to 5/10/20 shots in ChatGPT. Considering that different choices of few-shot training during training affect the test results, we repeatedly sample the same data on K random seeds simultaneously and calculate their mean values as reported. Considering more intuitive approaches to training the ChatGPT model, here are a few specific examples. As shown in Figure 1, examples are provided for training the model using few-shot (5-shot) methods, along with the corresponding results.
Table 4: The Accuracy, Precision, Recall, F1-scores results (%) on both English (DLC, CND and SCC) and Chinese (Sina, Tencent, Wechat and Paper) clickbait datasets in zero-shot scenario.
metrics Dataset
CND SCC DLC Sina Tencent Wechat Paper
Acc 63.66 53.87 76.44 61.23 53.80 62.02 63.59
Pre 47.88 53.69 75.87 54.61 53.13 57.71 61.71
Rec 74.79 49.68 77.53 57.84 64.40 90.00 71.54
F1 58.38 51.61 76.69 57.18 58.22 70.33 66.26
Table 5: The Accuracy, Precision, Recall, F1-scores results (%) on English datasets (DLC, CND and SCC). 200/5,300/10,and 600/20 denote the training number of not-prompt-tuning(CNN, BILSTM, GATED_CNN, FAST_TEXT and BERT) and ChatGPT methods. Ditto above.
Data Method Shot
200/5 300/10 600/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
CND CNN 77.86 79.36 77.77 79.27 80.50 79.00 80.83 79.12 80.59 80.70 80.64 80.35
BiLSTM 77.28 78.71 79.83 75.49 79.11 79.07 78.37 78.72 79.10 82.19 80.16 80.49
GATED_CNN 77.80 80.05 75.41 78.01 79.34 79.86 78.16 78.43 78.39 79.71 79.93 79.66
FAST_TEXT 65.23 72.31 63.84 72.32 74.89 71.07 74.30 65.58 75.13 78.33 73.60 77.63
BERT 71.25 72.00 71.25 71.00 77.54 78.28 76.46 77.49 82.52 83.16 81.08 82.94
ChatGPT 80.75 69.06 82.79 73.56 79.33 69.51 88.02 70.88 82.07 75.66 82.64 73.67
300/5 600/10 1200/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
SCC CNN 73.74 69.70 71.57 68.82 76.47 74.83 74.33 75.70 73.19 72.05 75.27 66.61
BiLSTM 75.18 75.78 74.06 73.50 75.74 79.54 78.28 79.15 71.79 81.70 80.08 82.71
GATED_CNN 70.92 71.41 70.18 67.69 71.26 74.53 76.04 72.28 68.00 74.54 74.23 75.39
FAST_TEXT 54.18 66.01 64.12 61.22 72.51 71.61 69.57 53.51 79.25 77.76 74.47 79.69
BERT 73.17 73.19 73.17 73.16 76.90 80.65 76.90 76.71 83.28 85.69 83.02 83.80
ChatGPT 60.00 59.02 68.36 62.06 59.44 60.76 61.00 60.13 60.90 63.33 66.88 62.57
500/5 1000/10 2000/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
DLC CNN 91.59 91.51 91.56 91.66 93.32 93.04 93.19 92.84 94.53 94.14 94.71 94.72
BiLSTM 86.69 86.45 85.91 86.48 87.36 86.08 88.20 87.97 89.58 88.45 88.24 89.14
GATED_CNN 92.07 91.78 91.90 91.67 93.56 93.14 93.27 93.56 93.22 94.67 94.28 94.45
FAST_TEXT 84.09 83.66 83.18 82.52 85.99 85.76 85.27 85.74 86.67 86.96 86.60 86.68
BERT 78.82 79.23 78.16 78.18 85.81 88.91 85.81 85.83 91.42 92.54 91.42 91.37
ChatGPT 83.94 88.63 90.01 83.08 86.14 92.41 86.93 85.09 88.73 95.54 87.81 87.82
Table 6: The Accuracy, Precision, Recall, F1-scores results (%) on Chinese datasets (Sina, Tencent, Wechat and Paper). 200/5,400/10,and 800/20 denote the training number of not-prompt-tuning(CNN, BILSTM, GATED_CNN, FAST_TEXT and BERT) and ChatGPT methods. Ditto above.
Data Method Shot
200/5 400/10 800/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Sina CNN 72.13 72.02 72.47 72.70 77.00 74.85 75.61 75.84 75.96 77.60 77.35 78.19
BiLSTM 74.74 71.83 58.71 70.88 74.39 72.32 72.65 75.40 75.60 76.11 72.65 74.01
GATED_CNN 73.87 72.09 71.25 71.37 75.61 76.07 77.35 75.66 77.00 76.20 75.61 76.93
FAST_TEXT 56.10 62.23 52.79 63.94 74.04 71.45 72.13 72.74 73.17 73.76 74.04 71.81
BERT 52.36 52.92 44.25 48.20 55.25 57.69 46.51 51.50 65.96 71.43 55.56 62.50
ChatGPT 70.73 76.21 60.28 67.32 69.16 71.65 63.41 67.28 75.30 83.80 62.85 71.83
400/5 800/10 1600/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Tencent CNN 54.58 53.97 56.94 54.42 59.55 56.16 59.16 57.41 61.65 59.29 59.69 62.75
BiLSTM 53.80 54.06 50.65 52.87 53.27 58.43 59.69 54.90 59.03 59.87 57.85 57.56
GATED_CNN 55.37 56.28 56.54 55.63 60.60 60.29 57.33 55.71 60.08 60.74 57.98 60.31
FAST_TEXT 53.53 50.58 51.05 56.49 56.28 57.75 57.07 49.95 55.37 57.94 58.12 58.20
BERT 47.71 47.55 43.19 45.27 52.80 54.20 47.26 50.49 59.54 62.00 53.26 57.30
ChatGPT 59.34 57.91 71.13 63.84 56.02 54.71 69.90 61.38 59.08 57.17 71.99 63.73
400/5 800/10 1600/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Wechat CNN 72.14 70.64 69.40 70.60 86.67 84.38 85.12 83.59 85.83 86.70 87.38 86.72
BiLSTM 64.29 61.93 67.74 62.24 81.67 81.52 76.55 81.61 83.57 81.78 80.95 83.49
GATED_CNN 67.74 69.17 69.17 67.62 83.45 86.31 84.29 85.27 87.74 88.16 86.67 86.02
FAST_TEXT 58.10 62.59 60.48 61.56 81.07 83.26 81.79 83.13 84.17 84.60 84.29 81.18
BERT 45.05 44.76 41.67 43.16 49.54 50.73 44.73 47.54 57.06 59.39 51.48 55.15
ChatGPT 60.12 56.69 85.71 68.24 60.41 56.16 93.99 70.26 64.48 59.93 90.24 71.01
800/5 1600/10 3200/20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Paper CNN 68.87 73.27 68.76 66.11 71.45 74.28 70.47 71.16 72.38 72.12 71.32 72.29
BiLSTM 64.36 68.41 65.96 65.89 68.44 68.07 66.26 64.94 69.78 68.23 69.03 69.85
GATED_CNN 68.44 72.40 68.63 67.29 71.41 74.78 69.36 65.84 71.79 72.02 71.04 72.91
FAST_TEXT 67.93 70.25 66.03 65.53 68.45 72.88 68.96 67.85 71.10 72.15 72.23 70.39
BERT 50.16 50.20 40.02 44.54 51.92 52.21 41.51 46.26 54.17 54.48 43.60 48.44
ChatGPT 68.33 72.18 62.70 66.42 68.32 70.40 63.88 66.98 70.94 72.14 67.21 69.59
Implementation details.
We chose the latest available LLMs GPT3.5 (gpt-3.5-turbo). In order to interact with the model, we access it by making API calls to the ChatGPT API, in which calling GPT3.5 requires payment. This model’s maximum context length is 4096 tokens, approximating 100 news headlines. Moreover, this model can only call three news headlines a minute. In addition, for the few-shot experiments, we select a few sentences that ChatGPT got wrong from the sevendata sets mentioned above as examples of clickbait detection. Figure 1 shows ChatGPT detecting clickbait errors.
Table 7: Comparison of different prompts for GPT3.5 on English datasets (%).
Data Method Shot
5 10 20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
CND GPT3.5+P1 80.75 69.06 78.69 73.56 79.33 69.51 70.08 69.79 82.07 75.66 70.08 72.76
GPT3.5+P2 67.98 52.20 82.79 64.03 75.25 59.33 88.02 70.88 79.80 66.45 82.64 73.67
SCC GPT3.5+P1 58.41 59.02 49.36 53.76 59.44 60.76 50.08 54.91 60.90 63.33 49.20 55.38
GPT3.5+P2 60.00 56.83 68.36 62.06 59.00 59.28 61.00 60.13 60.63 58.78 66.88 62.57
DLC GPT3.5+P1 83.94 88.63 78.19 83.08 86.14 92.41 78.85 85.09 88.73 95.54 81.26 87.82
GPT3.5+P2 81.49 76.54 90.01 82.73 84.39 82.71 86.93 84.77 87.52 87.30 87.81 87.55
Table 8: Comparison of different prompts for GPT3.5 on Chinses datasets (%).
Data Method Shot
5 10 20
Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
Sina GPT3.5+P1 65.33 68.80 56.10 61.80 60.80 60.69 61.32 61.00 63.94 64.60 61.67 63.10
GPT3.5+P2 70.73 76.21 60.28 67.32 69.16 71.65 63.41 67.28 75.30 83.80 62.85 0.7183
Tencent GPT3.5+P1 59.34 57.91 71.13 63.84 52.88 52.89 52.62 52.75 58.51 57.17 67.80 62.03
GPT3.5+P2 58.64 56.96 70.68 63.08 56.02 54.71 69.99 61.38 59.08 57.17 71.99 63.73
Wechat GPT3.5+P1 60.12 56.69 85.71 68.24 60.41 56.10 93.99 70.26 58.57 55.25 90.24 68.54
GPT3.5+P2 55.71 53.99 77.38 63.60 58.69 56.16 79.29 65.75 64.48 59.93 87.11 71.01
Paper GPT3.5+P1 68.33 70.60 62.70 66.42 68.32 70.40 63.88 66.98 68.70 70.84 61.94 66.09
GPT3.5+P2 68.13 72.18 59.59 65.28 66.43 69.54 55.77 61.9 70.94 72.14 67.21 69.59
Evaluation metrics
To test the effect of detection, we conduct four evaluation metrics to evaluate our method in experiments, such as accuracy, precision, recall and F1-score.
accuracy The accuracy can be defined as the ratio of correctly predicted samples to the total number of samples.
$Acc=\frac{TP+TN}{TP+TN+FP+FN}$
precision The positive prediction rate can be defined as the ratio of correctly predicted positive samples to the total number of positive samples in the prediction.
$P=\frac{TP}{TP+FP}$
recall The positive precision can be defined as the ratio of correctly predicted positive samples to the total number of samples labeled positive.
$R=\frac{TP}{TP+FN}$
F1-score The F1-score can be defined as the harmonic average of precision and recall.
$F1=\frac{{2}\times{P}\times{R}}{P+R}$
4.4 Experimental Results
The main results of the experiments on clickbait detection in English and Chinese datasets on zero-shot and few-shot scenarios are listed in Table 4, 5 and 6 respectively. We can see that the performance of ChatGPT has not demonstrated the best results on all four metrics over the seven datasets. Compared to the deep neural networks and fine-tuning PLMs, ChatGPT has a lot of room for improvement in clickbait detection. Moreover, with the increasing of pre-trained information, the effectiveness of clickbait detection, sometimes, is worse. Therefore, we can observe that a small amount of pre-training information does not significantly impact the performance of ChatGPT in detecting clickbait. Moreover, we can observe that ChatGPT can achieve stable results on almost all evaluation metrics. The results confirmed that LLMs can be adapted to other languages, and ChatGPT is a monolithic model capable of supporting multiple languages, which makes it a comprehensive multilingual clickbait detection technique.
4.5 Ablation Study
We compared the results between different prompts (as shown in Table 1 and Table 2) for clickbait detection tasks. As same as the main experiments, we select GPT3.5 as a backbone for the prompts experiments. The results comparison of different Prompts is shown in Table 7 and Table 8.
Generally speaking, there is no significant performance gap between the first prompt (P1) and the second prompt (P2) in terms of effectiveness for few-shot clickbait detection. Specifically, the performance of P1 is significantly better than that of P2 in the dataset Wechat. In dataset Sina, the opposite is true.
5 Conclusion and Future Work
In this paper, we present a study of the performance of LLMs (ChatGPT with GPT3.5) for clickbait detection. During the benchmark experiments on both English and Chinese datasets, LLMs did not perform as well as the current state-of-the-art deep neural networks and fine-tuning PLMs methods in the realm of multilingual clickbait detection. In our subsequent efforts, we will try to design more effective methods with the help of LLMs that can significantly and consistently outperform SOTA clickbait detection methods.

### v4

@@
 Clickbait, which aims to induce users with some surprising and even thrilling headlines for increasing click-through rates, permeates almost all online content publishers, such as news portals and social media.
-Recently, Large Language Models (LLMs) have emerged as a powerful instrument and achieved tremendous success in a serious of NLP downstream tasks.
+Recently, Large Language Models (LLMs) have emerged as a powerful instrument and achieved tremendous success in a series of NLP downstream tasks.
 However, it is not yet known whether LLMs can be served as a high-quality clickbait detection system.
-In this paper, we analyze the performance of LLMs in the few-shot scenarios on a number of English and Chinese benchmark datasets.
+In this paper, we analyze the performance of LLMs in the few-shot and zero-shot scenarios on several English and Chinese benchmark datasets.
 Experimental results show that LLMs cannot achieve the best results compared to the state-of-the-art deep and fine-tuning PLMs methods.
-Different from the human intuition, the experiments demonstrated that LLMs cannot make satisfied clickbait detection just by the headlines.
+Different from human intuition, the experiments demonstrated that LLMs cannot make satisfied clickbait detection just by the headlines.
 1 Introduction
@@
 (2015).
-Clickbait refers to deliberately enticed users to click with some curious and chilling headlines, which are always unrelated to the real content or even the advertising promotion Chakraborty et al.
+Clickbait refers to deliberately enticing users to click with some curious and chilling headlines, which are always unrelated to the real content or even the advertising promotion Chakraborty et al.
 (2016).
@@
 (2022).
-In recent years, the research methods for clickbait detection evolved from feature engineering to neural networks and, more recently, into pre-trained language models.
-Feature engineering methods extracted features such as semantic and linguistic features for detection tasks Biyani et al.
+In the last decade, the research methods for clickbait detection evolved from feature engineering to neural networks and, more recently, into Pre-trained Language Models(PLMs).
+Feature engineering methods extracted features such as semantic and linguistic features for detection Biyani et al.
 (2016); Wei and Wan (2017).
-The deep neural networks methods can learn more abstract and higher-level features by disentangling explanatory factors of variations behind news title and content for clickbait detection Yoon et al.
+The deep neural network methods can learn more abstract and higher-level features by disentangling explanatory factors of variations behind news titles and content for clickbait detection Yoon et al.
 (2019); Zheng et al.
 (2021).
-In recent years, Pre-trained Language Models (PLMs), such as BERT Devlin et al.
-(2018), also show superiority in clickbait detection task.
-However, the feature engineering-based methods and deep neural networks typically require large-scaled labeled data since the detection is regarded as a classification method.
-In the PLMs methods, the huge gap between pre-training and fine-tuning prevent detection tasks from fully utilizing pre-training knowledge.
+Recently, fine-tuning PLMs methods, such as BERT Devlin et al.
+(2018), has shown superiority in clickbait detection tasks.
+However, feature engineering-based methods and deep neural networks typically require large-scale labeled data since the detection is regarded as a classification method.
+In the PLMs methods, the huge gap between pre-training and fine-tuning prevents detection tasks from fully utilizing pre-training knowledge.
 More recently, Large Language Models (LLMs) have demonstrated the powerful ability in various NLP downstream tasks Brown et al.
@@
 (2022); Thoppilan et al.
-(2022), which can achieve awesome performance even in the few-shot and zero-shot scenarios.
+(2022), which can achieve awesome performance even in the few-shot or even zero-shot scenarios.
 Nevertheless, it remains unclear how LLMs perform in clickbait detection tasks compared to current methods.
-To address this issue, in this paper, we conduct a systematic evaluation of the few-shot and zero-shot learning capability of LLMs.
-The experiments are validated on both English and Chinese open datasets, and we conducted an empirical comparison of the performance between ChatGPT with the GPT3.5 model (gpt-3.5-turbo) and other state-of-the-art methods such as prompt-tuning.
-To the best of our knowledge, this is the first attempt to validate the performance of LLMs on clickbait detection, the works in this paper provide a preliminary evaluation including detection results and robustness.
+To address this issue, in this paper, we conduct a systematic evaluation of the few-shot and zero-shot learning capabilities of LLMs.
+The experiments were validated on English and Chinese open datasets, and we conducted a performance comparison between ChatGPT with the GPT-3.5 model (gpt-3.5-turbo) and GPT-4 model, and other state-of-the-art methods such as prompt-tuning.
+To the best of our knowledge, this is the first attempt to validate the performance of LLMs on clickbait detection, we provide a preliminary evaluation including detection results and robustness.
 The key findings and insights are summarized as follows:
-  • ChatGPT achieves unsatisfying results in the few-shot scenarios compared to the state-of-the-art clickbait detection methods.
+  • LLMs achieves unsatisfying results in the few-shot and zero-shot scenarios compared to the state-of-the-art clickbait detection methods.
 We found that fine-tuning PLMs and prompt-tuning can achieve better results just based on the titles, which is consistent with humans in the real-world.
@@
 This also confirms that LLMs can be adapted to other languages.
-It is worth mentioning that the source code and all the results of this paper is available on https://github.com/zhuyiYZU/chatGPTforClickbait.
+It is worth mentioning that the source code and all the results of this paper are available at https://github.com/zhuyiYZU/chatGPTforClickbait.
 2 Related Work
 2.1 Clickbait Detection
-Clickbait detection is an emerging field which is attracting increasing attention in recent years.
-On most online services, such as e-commerce, social media, and news portals, more clicks means more profit and commercial revenue.
+Clickbait detection is an emerging field that has attracted increasing attention in recent years.
+On most online services, such as e-commerce, social media, and news portals, more clicks mean more profit and commercial revenue.
 Early clickbait detection methods mainly focus on extracting a variety of features for detection tasks, such as semantic Rony et al.
@@
 2.2 Large Language Models
-Represented by GPT-3 Brown et al.
+Represented by GPT-4 Brown et al.
 (2020), Large Language Models (LLMs) have achieved superior performance, especially in few-shot learning scenarios Brown et al.
@@
 The first is the larger scale, LLMs have a much larger scale in terms of model parameters and training data.
-Secondly, without the fine-tuning PLMs, LLMs can prompt few-shot learning that requires no additional neural layer and shows excellent performance.
+Secondly, without fine-tuning PLMs, LLMs can prompt few-shot learning that requires no additional neural layer and shows excellent performance.
 However, there is no work on the capabilities of LLMs on clickbait detection tasks.
 3 Methodology
-Through the few-shot learning ability of LLMs, they can also achieve excellent performance for tasks with low resources.
-Considering the lack of large-scale training corpus for clickbait detection tasks, we will test the performance of LLMs in few-shot scenarios on clickbait detection.
+Through the rich knowledge in upstream large models, they can achieve excellent performance in downstream tasks with low resources.
+Considering the lack of large-scale training corpus for clickbait detection tasks, we will test the performance of LLMs in few-shot and zero-shot scenarios on clickbait detection.
 LLMs typically use prompts (i.e.
 specific templates) to guide the model in predicting output or answers, without requiring specific training on the data.
-Utilizing this form of prompt, we conducted experiments with different prompts on OpenAI’s largest available model GPT3.5 (gpt-3.5-turbo) for ChatGPT.
-Prompt.
+Utilizing this form of prompt, we conducted experiments with different prompts on OpenAI’s most representative available models GPT3.5 (gpt-3.5-turbo) and GPT-4 for ChatGPT.
+3.1 Prompt
 To stimulate the rich knowledge distributed in LLMs, we manually designed the prompts for validating the performance of clickbait detection.
@@
 For the two prompts, {Clickbait Sentence} and {Not-Clickbait Sentence} refer to the input sentence, {Yes, it is a clickbait} and {No, it is not a clickbait} are the corresponding labels respectively.
-{Results} is the place the carries the outputs of LLMs.
+{Results} is that place that carries the outputs of LLMs.
 Table 1: The hand-crafted prompts for clickbait detection.
-Prompts
-P1
-I want you to detect whether
-the input sentence is clickbait or not.
+P_id Prompts
+P1 I want you to detect whether the input sentence is clickbait or not.
 Input: {Clickbait Sentence}
@@
 In the first prompt (P1), the {guide-input-output} pattern is employed to guide LLMs for detecting clickbaits.
-In the second prompt (P2), the {sentence-question-answer} pattern is utilized to detect clickbaits in the form of a question.
-It is worth mentioning that the specialized guide word, such as ”Output:” and ”Answer:”, are added to the prompts, then the clickbait detection is regarded as the close-style task for LLMs, which ensured to achievement of a unique output.
-When performing multilingual clickbait detection tasks, we translate these two prompts into Chinese used in specific tasks as shown in Table 2.
-Table 2: The hand-crafted prompts for clickbait detection.
-提示
-P1
-请帮我检测
-以下输入是或不是点击诱饵。
-输入: {点击诱饵句子}
-输出:{是,这是一个点击诱饵}
-输入: {非点击诱饵句子}
-输出: {不,这不是一个点击诱饵}
-输入: {点击诱饵或非点击诱饵句子}
-输出: {结果}
-P2 句子:{点击诱饵句子}
-问题: 检测上面的句子是或不是点击诱饵。
-回答: {是,这是一个点击诱饵}
-句子:{非点击诱饵句子}
-问题: 检测上面的句子是或不是点击诱饵。
-回答: {不,这不是一个点击诱饵}
-句子:{点击诱饵或非点击诱饵句子}
-问题: 检测上面的句子是或不是点击诱饵。
-回答: {结果}
-Zero-shot. In the clickbait detection experiments in zero-shot scenarios, just one {guide-input-output} or {sentence-question-answer} pattern is directly input into the LLMs for the detection result.
-Few-shot.
-In the few-shot scenario, some instances follow the {guide-input-output} or {sentence-question-answer} pattern are provided as the training data for the detection tasks.
-Notably, in the first prompt (P1), the guidance is not needed to be repeated, only the input-output is repeated for some instance sentences and corresponding labels are stacked.
-In the second prompt (P2), the sentence-question-answer is repeated for each sample as training data.
-(a) English P1
-(b) Chinese P1
-(c) English P2
-(d) Chinese P2
-Figure 1: The example of ChatGPT detecting clickbait errors.
-4 Experiment
-4.1 Datasets and Templates
-To evaluate the performance of our method for clickbait detection in both English and Chinese, we conduct experiments on seven datasets.
-The DL-Clickbait (DLC), Clickbait news detection (CND), SC-Clickbait (SCC) are three well-known public clickbait detection datasets in English.
-Sina, Tencent, Wechat, and Paper are four public clickbait detection datasets in Chinese.
+In the second prompt (P2), the {sentence-question-answer} pattern is utilized to detect clickbait in the form of a question.
+It is worth mentioning that the specialized guide words, such as ”Output:” and ”Answer:”, are added to the prompts, then the clickbait detection is regarded as the close-style task for LLMs, which ensures to achievement of a unique output.
+When performing multilingual clickbait detection tasks, we translate these two prompts into Chinese used in specific tasks .
+4 Experiments
+4.1 Datasets
+To evaluate the performance of LLM for clickbait detection in English and Chinese, we conducted experiments on seven datasets.
+The DL-Clickbait (DLC), Clickbait news detection (CND), and SC-Clickbait (SCC) are three well-known public English datasets.
+Sina, Tencent, WeChat, and Paper are four Chinese clickbait detection datasets.
 The statistical details of all the datasets are presented in Table 2.
@@
 Statistic DLC CND SCC Sina Tencent WeChat Paper
-#total headline 2345 8324 32000 1912 2546 2800 30800
-#train sets 1638 5801 22400 1338 1782 1960 21560
-#test sets 707 2523 9600 574 764 840 9240
-avg.# words per sentence 11 13 9 21 25 21 22
-4.2 Baselines
-We compare ChatGPT with the following deep neural network and fine-tuning PLMs methods.
-TextCNN Agrawal (2016) TextCNN utilizes kernels with different sizes to find various features from the input text.
+#total headline 2388 8324 32000 1648 2594 2808 30800
+#train sets 1671 5801 22400 1153 1815 1965 21560
+#test sets 717 2523 9600 495 779 843 9240
+avg.# words per sentence 67 13 53 604 316 195 621
+4.2 Baseline
+We compare our ChatGPT with the following deep neural network and fine-tuning PLMs methods.
+CNN Agrawal (2016) CNN utilizes kernels with different sizes to find various features from the input text.
 These features are automatically detected and are used to train a neural network for the corresponding task.
@@
 It used a Siamese net to detect similarities between the source and target data.
-To add another layer of complexity to the model, it also use TextCNN to learn image embeddings from large amounts of data.
 GATED_CNN Yang et al. (2019) The method can be stacked to represent large context sizes and extract hierarchical features over larger and larger contexts with more abstractive features.
@@
 BERT Devlin et al. (2018) The BERT method uses words and sentences to distinguish the context of words in a sentence.
+MFWCD-BERT Liu et al.
+(2022) The multi-feature WeChat clickbait detection method, which incorporates semantic, syntactic, and auxiliary information, achieves excellent performance.
+It leverages an enhanced Graph Attention Network (GAT) to aggregate title and local syntactic structures, using an attention mechanism to capture valuable structural features.
+PT Liu et al.
+(2023) PT is a technique known as prompt tuning, which incorporates a cloze-style task during the model fine-tuning process.
+This approach enables a more effective utilization of pre-training information, thereby further enhancing the model’s performance in few-shot learning scenarios.
+PEPT Wu et al.
+(2023) PEPT is a method that introduces prompt learning into clickbait detection.
+It effectively performs classification tasks and part-of-speech tagging using pre-trained language models by training them on specific tasks.
+It’s worth noting that this approach is inspired by typical grammar and semantic features of clickbait, and it achieves grammar-guided semantic understanding, which can perform well in both low-resource and full-scale scenarios.
 4.3 Implementation Details and Evaluation Metrics
-Few-shot settings
-In this experiment, we randomly selected K(5,10,20) instances as the training set for ChatGPT, and part of the samples are selected as the test set.
-Considering that too few training samples may greatly affect the effectiveness of the baselines, we draw different training samples from different datasets, where the number of training samples is 200/300/600, 300/600/1200, and 500/1000/2000 for English datasets (DL-Clickbait, CND, and SC-Clickbait datasets) and 200/400/800, 400/800/1600, 400/800/1600 and 800/1600/3200 for Chinese datasets (Sina, Tencent, Wechat and Paper), corresponding to 5/10/20 shots in ChatGPT.
+4.3.1 Few-shot settings
+In the experiments, we randomly selected K (5,10,20) instances as the training set for our ChatGPT, and selected part of the samples as the test set.
+Considering that too few training samples may greatly affect the effectiveness of the baselines, we select different training samples from different datasets, where the number of training samples is 300/600/1200, 400/800/1600, and 1000/2000/4000 for English datasets (DL-Clickbait, CND, and SC-Clickbait datasets) and 200/800/1600, 400/800/1600, 400/800/1600 and 800/1600/3200 for Chinese datasets(Sina, Tencent, WeChat and Paper), corresponding to 5/10/20 shots.
 Considering that different choices of few-shot training during training affect the test results, we repeatedly sample the same data on K random seeds simultaneously and calculate their mean values as reported.
@@
 As shown in Figure 1, examples are provided for training the model using few-shot (5-shot) methods, along with the corresponding results.
-Table 4: The Accuracy, Precision, Recall, F1-scores results (%) on both English (DLC, CND and SCC) and Chinese (Sina, Tencent, Wechat and Paper) clickbait datasets in zero-shot scenario.
-metrics Dataset
-CND SCC DLC Sina Tencent Wechat Paper
-Acc 63.66 53.87 76.44 61.23 53.80 62.02 63.59
-Pre 47.88 53.69 75.87 54.61 53.13 57.71 61.71
-Rec 74.79 49.68 77.53 57.84 64.40 90.00 71.54
-F1 58.38 51.61 76.69 57.18 58.22 70.33 66.26
-Table 5: The Accuracy, Precision, Recall, F1-scores results (%) on English datasets (DLC, CND and SCC).
-200/5,300/10,and 600/20 denote the training number of not-prompt-tuning(CNN, BILSTM, GATED_CNN, FAST_TEXT and BERT) and ChatGPT methods.
+(a) English P1
+(b) Chinese P1
+(c) English P2
+(d) Chinese P2
+Figure 1: The example of GPT-3.5 detecting clickbait errors.
+4.3.2 Parameter settings
+For using ChatGPT for clickbait detection, we have chosen the most representative available models GPT3.5 (gpt-3.5-turbo) and GPT-4, to interact with.
+To establish communication with the model, we have opted to use the ChatGPT API.
+First, we need to register and apply for a ChatGPT API key from OpenAI.
+This API key will be used to access the API.
+Then, we send an HTTP POST request to the API endpoint to engage in a conversation with the ChatGPT model.In this process, we set up an array of messages, including user messages and system messages, corresponding to the two roles, ”user” and ”assistant”.
+Based on the input of users, messages will be alternately added to the array as user messages and system messages.
+Subsequently, we extract the assistant’s responses from the JSON response returned by the API.It’s important to note that using the ChatGPT API incurs a cost, and the text content of API requests must not exceed the maximum token limit of the model, which, in this case, is 4096 tokens, roughly equivalent to 70 news headlines.
+Additionally, there are rate limits for API requests, with a maximum of three news headlines per minute and no more than two hundred requests in a day.In addition, for the few-shot experiments, we select a few sentences that GPT-3.5 got wrong from the seven data sets mentioned above as examples of clickbait detection.
+Figure 1 shows ChatGPT detecting clickbait errors.
+We also provide a detailed description of the parameter settings for our text classification models (CNN, BiLSTM, and FastText) used in the clickbait detection task.
+For the CNN model, we employ three convolutional layers, each with a kernel size of 3x3 and 64, 128, and 256 kernels, respectively.
+The stride is set to 1, and padding is ”same.” Following each convolutional layer, we incorporate max-pooling layers with a kernel size of 2x2 and a stride of 2.
+After the convolutional layers, we have two fully connected layers with 512 and 256 neurons, respectively.
+Additionally, we apply ReLU activation functions after each convolutional and fully connected layer.
+For the BiLSTM model, we utilize two layers of bidirectional LSTM, with 64 and 128 LSTM units in each layer.
+Since the BiLSTM layers are bidirectional, the total number of LSTM units is 256 (128 * 2).
+After the BiLSTM layers, we add two fully connected layers with 512 and 256 neurons, respectively.
+Similarly, ReLU activation functions are applied after each fully connected layer.
+In the case of the FastText model, we represent the text data as bag-of-words vectors.
+We set up a fully connected layer with two neurons, corresponding to the binary classification task.
+After the fully connected layer, we use the Sigmoid activation function.
+For all three models, we employ binary cross-entropy loss functions as the objective functions to measure model performance.
+We use the Adam optimizer, with a typical training batch size of 64 and 50 training epochs.
+We conducted hyperparameter tuning, including learning rates, the number of convolutional kernels, LSTM unit numbers, bag-of-words sizes, and word vector dimensions, selecting the best hyperparameter values through random search.
+For the BERT model, we employed the BERT-base-uncased variant and conducted training over 10 epochs using a batch size of 32.
+For training prompt-based learning models (PT, PEPT), we utilized the xlm-roberta-large pretrained language model.
+To prevent overfitting, we applied a dropout rate of 0.5 during the training process.
+We configured the learning rate at 3e-5 and set the batch size to 16, incorporating a weight decay of 1e-5.
+Additionally, we performed validation steps to fine-tune the model’s hyperparameters.
+Across all our small-scale experiments, we maintained a consistent number of 10 epochs for comprehensive training and reliable results.
+We employed the Adam optimizer to optimize the model parameters.
+The experimental outcomes were derived from a server configuration featuring an NVIDIA GeForce RTX 3090 Founders Edition GPU, an Intel(R) Core(TM) i9-10980XE CPU clocked at 3.00 GHz, and 125 GB of RAM.
+4.3.3 Evaluation metrics
+To test the effect of detection, we conduct four evaluation metrics to evaluate our method in experiments, such as accuracy, precision, recall and F1-score.
+accuracy The accuracy can be defined as the ratio of correctly predicted samples to the total number of samples.
+$Acc=\frac{TP+TN}{TP+TN+FP+FN}$ (1)
+precision The positive prediction rate can be defined as the ratio of correctly predicted positive samples to the total number of positive samples in the prediction.
+$Pre=\frac{TP}{TP+FP}$ (2)
+recall The positive precision can be defined as the ratio of correctly predicted positive samples to the total number of samples labeled positive.
+$Rec=\frac{TP}{TP+FN}$ (3)
+F1-score The F1-score can be defined as the harmonic average of precision and recall.
+$F1=\frac{{2}\times{Pre}\times{Rec}}{Pre+Rec}$ (4)
+Table 3: The Accuracy, Precision, Recall, F1-scores results (%) on English datasets ( DLC, CND, and SCC).
+200/5,300/10, and 600/20 denote the training number of not-prompt-tuning(CNN, BILSTM, GATED-CNN, FAST-TEXT, BERT, and MFWCD-BERT) and prompt-tuning(PT, PEPL, GPT-3.5 and GPT-4) methods.
 Ditto above.
+The bolder ones mean better.
 Data Method Shot
-200/5 300/10 600/20
-Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
-CND CNN 77.86 79.36 77.77 79.27 80.50 79.00 80.83 79.12 80.59 80.70 80.64 80.35
-BiLSTM 77.28 78.71 79.83 75.49 79.11 79.07 78.37 78.72 79.10 82.19 80.16 80.49
-GATED_CNN 77.80 80.05 75.41 78.01 79.34 79.86 78.16 78.43 78.39 79.71 79.93 79.66
-FAST_TEXT 65.23 72.31 63.84 72.32 74.89 71.07 74.30 65.58 75.13 78.33 73.60 77.63
-BERT 71.25 72.00 71.25 71.00 77.54 78.28 76.46 77.49 82.52 83.16 81.08 82.94
-ChatGPT 80.75 69.06 82.79 73.56 79.33 69.51 88.02 70.88 82.07 75.66 82.64 73.67
 300/5 600/10 1200/20
 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
-SCC CNN 73.74 69.70 71.57 68.82 76.47 74.83 74.33 75.70 73.19 72.05 75.27 66.61
+DLC CNN 60.39 70.95 60.39 65.24 70.29 68.71 70.29 69.49 81.72 81.55 81.72 81.63
+BiLSTM 73.08 72.88 73.08 72.97 77.68 77.87 77.68 77.77 82.28 82.89 82.28 82.58
+GATED-CNN 77.80 80.05 75.41 78.01 79.34 79.86 78.16 78.43 78.39 79.71 79.93 79.66
+FAST-TEXT 65.23 72.31 63.84 72.32 74.89 71.07 74.30 65.58 75.13 78.33 73.60 77.63
+BERT 71.86 71.83 71.86 71.84 77.40 80.62 77.40 78.97 82.42 82.13 82.42 82.27
+MFWCD-BERT 79.13 80.44 79.13 78.91 82.19 82.21 82.19 82.19 85.00 85.90 85.00 84.91
+PT 81.86 82.69 81.86 80.60 82.70 82.43 82.70 82.42 84.79 84.60 84.79 84.52
+PEPL 83.31 82.70 79.29 80.53 87.41 90.11 82.40 84.74 88.97 89.58 85.67 87.16
+[…]
+GPT-4 90.54 94.00 77.05 84.68 89.96 96.24 73.36 83.26 86.05 88.3 68.03 76.85
+300/5 600/10 1200/20
+Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
+CND CNN 73.74 69.70 71.57 68.82 76.47 74.83 74.33 75.70 73.19 72.05 75.27 66.61
 BiLSTM 75.18 75.78 74.06 73.50 75.74 79.54 78.28 79.15 71.79 81.70 80.08 82.71
-GATED_CNN 70.92 71.41 70.18 67.69 71.26 74.53 76.04 72.28 68.00 74.54 74.23 75.39
-FAST_TEXT 54.18 66.01 64.12 61.22 72.51 71.61 69.57 53.51 79.25 77.76 74.47 79.69
+GATED-CNN 70.92 71.41 70.18 67.69 71.26 74.53 76.04 72.28 68.00 74.54 74.23 75.39
+FAST-TEXT 54.18 66.01 64.12 61.22 72.51 71.61 69.57 53.51 79.25 77.76 74.47 79.69
 BERT 73.17 73.19 73.17 73.16 76.90 80.65 76.90 76.71 83.28 85.69 83.02 83.80
-ChatGPT 60.00 59.02 68.36 62.06 59.44 60.76 61.00 60.13 60.90 63.33 66.88 62.57
-500/5 1000/10 2000/20
+MFWCD-BERT 75.07 75.18 74.98 74.99 78.09 79.54 79.11 78.97 80.58 81.89 80.16 80.94
+PT 53.07 54.04 53.07 50.05 62.89 63.34 62.89 62.57 66.58 68.09 66.58 65.86
+PEPL 58.19 61.50 58.19 54.95 62.73 64.47 62.73 61.58 63.08 63.46 63.08 62.83
+[…]
+GPT-4 68.61 69.47 66.97 68.2 64.32 64.65 63.2 63.92 63.47 72.56 43.32 54.25
+1000/5 2000/10 4000/20
 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
-DLC CNN 91.59 91.51 91.56 91.66 93.32 93.04 93.19 92.84 94.53 94.14 94.71 94.72
-BiLSTM 86.69 86.45 85.91 86.48 87.36 86.08 88.20 87.97 89.58 88.45 88.24 89.14
-GATED_CNN 92.07 91.78 91.90 91.67 93.56 93.14 93.27 93.56 93.22 94.67 94.28 94.45
-FAST_TEXT 84.09 83.66 83.18 82.52 85.99 85.76 85.27 85.74 86.67 86.96 86.60 86.68
-BERT 78.82 79.23 78.16 78.18 85.81 88.91 85.81 85.83 91.42 92.54 91.42 91.37
-ChatGPT 83.94 88.63 90.01 83.08 86.14 92.41 86.93 85.09 88.73 95.54 87.81 87.82
-Table 6: The Accuracy, Precision, Recall, F1-scores results (%) on Chinese datasets (Sina, Tencent, Wechat and Paper).
-200/5,400/10,and 800/20 denote the training number of not-prompt-tuning(CNN, BILSTM, GATED_CNN, FAST_TEXT and BERT) and ChatGPT methods.
+SCC CNN 75.31 76.30 75.31 75.80 82.28 82.01 82.28 82.14 86.19 86.11 86.19 86.14
+BiLSTM 79.77 79.36 79.77 79.55 83.22 83.54 83.22 83.37 87.09 87.86 87.09 87.47
+GATED-CNN 92.07 91.78 91.90 91.67 93.56 93.14 93.27 93.56 93.22 94.67 94.28 94.45
+FAST-TEXT 84.09 83.66 83.18 82.52 85.99 85.76 85.27 85.74 86.67 86.96 86.60 86.68
+BERT 78.52 78.04 78.52 78.27 83.96 83.78 83.96 83.86 89.92 89.95 89.92 89.93
+MFWCD-BERT 81.94 81.95 81.94 81.94 83.82 83.91 83.82 83.81 85.69 86.10 85.69 85.65
+PT 90.91 91.12 90.91 90.90 92.94 92.45 92.47 92.42 93.37 93.44 93.37 93.37
+PEPL 83.10 85.57 83.11 82.81 95.00 95.00 95.00 95.00 96.53 96.54 96.53 96.53
+[…]
+GPT-4 94.62 98.83 90.3 94.37 96.88 97.92 95.80 96.85 97.17 98.79 95.5 97.12
+Table 4: The Accuracy, Precision, Recall, F1-scores results (%) on Chinese datasets (Sina, Tencent, WeChat, and Paper).
+200/5,400/10, and 800/20 denote the training number of not-prompt-tuning(CNN, BILSTM, GATED-CNN, FAST-TEXT, BERT, and MFWCD-BERT) and prompt-tuning(PT, PEPL, GPT-3.5 and GPT-4) methods.
 Ditto above.
+The bolder ones mean better.
 Data Method Shot
-200/5 400/10 800/20
+200/5 800/10 1600/20
 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
-Sina CNN 72.13 72.02 72.47 72.70 77.00 74.85 75.61 75.84 75.96 77.60 77.35 78.19
-BiLSTM 74.74 71.83 58.71 70.88 74.39 72.32 72.65 75.40 75.60 76.11 72.65 74.01
-GATED_CNN 73.87 72.09 71.25 71.37 75.61 76.07 77.35 75.66 77.00 76.20 75.61 76.93
-FAST_TEXT 56.10 62.23 52.79 63.94 74.04 71.45 72.13 72.74 73.17 73.76 74.04 71.81
-BERT 52.36 52.92 44.25 48.20 55.25 57.69 46.51 51.50 65.96 71.43 55.56 62.50
-ChatGPT 70.73 76.21 60.28 67.32 69.16 71.65 63.41 67.28 75.30 83.80 62.85 71.83
+Sina CNN 64.52 59.23 64.52 61.76 69.62 67.09 69.62 68.33 79.51 77.40 79.51 78.44
+BiLSTM 70.04 70.26 70.04 70.15 74.48 74.38 74.48 74.43 81.72 80.98 81.72 81.35
+GATED-CNN 73.87 72.09 71.25 71.37 75.61 76.07 77.35 75.66 77.00 76.20 75.61 76.93
+FAST-TEXT 56.10 62.23 52.79 63.94 74.04 71.45 72.13 72.74 73.17 73.76 74.04 71.81
+BERT 52.36 52.86 44.25 48.20 55.25 57.69 46.51 51.50 65.96 71.42 55.56 62.50
+MFWCD-BERT 77.80 79.49 77.41 77.24 82.43 85.15 82.93 82.39 85.28 88.56 85.22 85.77
+PT 79.14 79.17 79.13 78.91 82.54 82.65 82.58 82.11 86.05 86.12 86.17 86.08
+PEPL 77.51 80.04 77.24 77.34 79.28 79.17 79.13 79.20 84.92 85.44 84.62 84.01
+[…]
+GPT-4 73.87 93.08 51.57 66.37 78.36 91.33 62.59 74.27 75.61 82.67 64.81 72.66
 400/5 800/10 1600/20
 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
-Tencent CNN 54.58 53.97 56.94 54.42 59.55 56.16 59.16 57.41 61.65 59.29 59.69 62.75
-BiLSTM 53.80 54.06 50.65 52.87 53.27 58.43 59.69 54.90 59.03 59.87 57.85 57.56
-GATED_CNN 55.37 56.28 56.54 55.63 60.60 60.29 57.33 55.71 60.08 60.74 57.98 60.31
-FAST_TEXT 53.53 50.58 51.05 56.49 56.28 57.75 57.07 49.95 55.37 57.94 58.12 58.20
-BERT 47.71 47.55 43.19 45.27 52.80 54.20 47.26 50.49 59.54 62.00 53.26 57.30
-ChatGPT 59.34 57.91 71.13 63.84 56.02 54.71 69.90 61.38 59.08 57.17 71.99 63.73
+Tencent CNN 68.50 69.31 68.50 68.91 74.43 73.78 74.43 74.11 80.19 81.08 80.19 80.63
+BiLSTM 71.49 73.48 71.49 72.47 74.51 73.90 74.51 74.22 82.30 82.60 82.30 82.46
+GATED-CNN 55.37 56.28 56.54 55.63 60.60 60.29 57.33 55.71 60.08 60.74 57.98 60.31
+FAST-TEXT 53.53 50.58 51.05 56.49 56.28 57.75 57.07 49.95 55.37 57.94 58.12 58.20
+BERT 79.28 78.93 79.28 79.11 82.66 83.40 82.66 83.03 85.52 83.22 85.52 84.38
+MFWCD-BERT 81.44 86.34 81.40 81.02 83.78 84.01 83.70 83.41 87.71 88.03 87.61 87.20
+PT 79.72 82.48 79.82 79.46 81.76 84.48 81.35 78.62 82.59 82.65 82.58 82.54
+PEPL 74.87 74.90 74.82 74.70 76.72 77.03 76.65 76.61 80.03 80.05 80.00 80.81
+[…]
+GPT-4 55.97 67.76 25.94 37.52 58.32 66.16 34.29 45.17 60.99 60.14 65.18 62.56
 400/5 800/10 1600/20
 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
-Wechat CNN 72.14 70.64 69.40 70.60 86.67 84.38 85.12 83.59 85.83 86.70 87.38 86.72
-BiLSTM 64.29 61.93 67.74 62.24 81.67 81.52 76.55 81.61 83.57 81.78 80.95 83.49
-GATED_CNN 67.74 69.17 69.17 67.62 83.45 86.31 84.29 85.27 87.74 88.16 86.67 86.02
-FAST_TEXT 58.10 62.59 60.48 61.56 81.07 83.26 81.79 83.13 84.17 84.60 84.29 81.18
-BERT 45.05 44.76 41.67 43.16 49.54 50.73 44.73 47.54 57.06 59.39 51.48 55.15
-ChatGPT 60.12 56.69 85.71 68.24 60.41 56.16 93.99 70.26 64.48 59.93 90.24 71.01
-800/5 1600/10 3200/20
+WeChat CNN 77.08 78.70 77.08 77.88 82.05 80.34 82.05 81.19 83.35 81.02 83.35 82.17
+BiLSTM 78.86 77.41 78.86 78.13 83.80 83.53 83.80 83.67 85.20 85.78 85.20 85.49
+GATED-CNN 67.74 69.17 69.17 67.62 83.45 86.31 84.29 85.27 87.74 88.16 86.67 86.02
+FAST-TEXT 58.10 62.59 60.48 61.56 81.07 83.26 81.79 83.13 84.17 84.60 84.29 81.18
+BERT 83.19 79.38 83.19 81.24 85.11 84.95 85.11 85.03 88.14 89.00 88.14 88.57
+MFWCD-BERT 85.92 89.12 85.81 85.76 88.86 89.59 88.79 88.31 90.31 92.12 90.58 91.21
+PT 78.27 78.54 78.27 78.25 81.37 82.34 81.40 81.32 83.96 84.72 83.96 83.87
+PEPL 76.35 81.10 76.37 76.18 80.11 82.27 80.45 80.25 84.55 86.61 84.26 84.03
+[…]
+GPT-4 73.79 69.18 88.57 77.68 73.18 66.5 93.57 77.74 74.23 66.56 98.12 79.32
+400/5 800/10 1600/20
 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
-Paper CNN 68.87 73.27 68.76 66.11 71.45 74.28 70.47 71.16 72.38 72.12 71.32 72.29
-BiLSTM 64.36 68.41 65.96 65.89 68.44 68.07 66.26 64.94 69.78 68.23 69.03 69.85
-GATED_CNN 68.44 72.40 68.63 67.29 71.41 74.78 69.36 65.84 71.79 72.02 71.04 72.91
-FAST_TEXT 67.93 70.25 66.03 65.53 68.45 72.88 68.96 67.85 71.10 72.15 72.23 70.39
-BERT 50.16 50.20 40.02 44.54 51.92 52.21 41.51 46.26 54.17 54.48 43.60 48.44
-ChatGPT 68.33 72.18 62.70 66.42 68.32 70.40 63.88 66.98 70.94 72.14 67.21 69.59
-Implementation details.
-We chose the latest available LLMs GPT3.5 (gpt-3.5-turbo).
-In order to interact with the model, we access it by making API calls to the ChatGPT API, in which calling GPT3.5 requires payment.
-This model’s maximum context length is 4096 tokens, approximating 100 news headlines.
-Moreover, this model can only call three news headlines a minute.
-In addition, for the few-shot experiments, we select a few sentences that ChatGPT got wrong from the sevendata sets mentioned above as examples of clickbait detection.
-Figure 1 shows ChatGPT detecting clickbait errors.
-Table 7: Comparison of different prompts for GPT3.5 on English datasets (%).
+Paper CNN 72.21 75.57 72.21 73.85 76.20 74.09 76.20 75.13 79.33 77.48 79.33 78.41
+BiLSTM 74.69 78.10 74.69 76.35 77.28 79.06 77.28 78.16 80.40 80.36 80.40 80.38
+GATED-CNN 68.44 72.40 68.63 67.29 71.41 74.78 69.36 65.84 71.79 72.02 71.04 72.91
+FAST-TEXT 67.93 70.25 66.03 65.53 68.45 72.88 68.96 67.85 71.10 72.15 72.23 70.39
+BERT 75.79 76.99 75.79 76.38 79.25 77.63 79.25 78.42 82.65 82.71 82.65 82.68
+MFWCD-BERT 78.75 80.01 78.69 78.90 80.84 80.83 80.79 81.03 84.54 85.02 84.11 84.51
+PT 72.07 76.98 72.24 71.33 75.61 76.05 75.61 75.51 80.41 80.55 80.41 80.39
+PEPL 70.14 74.46 70.14 70.53 73.62 73.88 73.62 73.55 78.28 78.77 78.28 78.19
+[…]
+GPT-4 77.75 80.04 73.85 76.82 72.6 66.99 89.09 76.48 62.28 57.34 95.82 71.75
+4.4 Experimental Results
+The main results of the experiments on clickbait detection in English and Chinese datasets are listed in Tables 3 and 4 respectively, which presents the detection results of ChatGPT (including GPT-3.5 and GPT-4) in the few-shot scenario, along with a comparison to other state-of-the-art methods.
+We can see that the performance of ChatGPT has not demonstrated the best results on all four metrics over the seven datasets.
+Compared to the deep neural networks and fine-tuning PLMs, ChatGPT has a lot of room for improvement in clickbait detection, especially the performance of GPT-3.5 on the Chinese dataset.
+In addition, with the increase of pre-trained information, the effectiveness of clickbait detection, sometimes, is worse.
+Therefore, we can observe that a small amount of pre-training information does not significantly impact the performance of ChatGPT in detecting clickbait.
+Moreover, we can observe that ChatGPT can achieve stable results on almost all evaluation metrics.
+The results confirmed that LLMs can be adapted to other languages, and ChatGPT is a monolithic model capable of supporting multiple languages, which makes it a comprehensive multilingual clickbait detection technique.
+It is worth noting that GPT-4 shows significant improvements over GPT-3.5 in areas such as comprehension and handling complex tasks, especially in scenarios that require a high level of text understanding, such as clickbait detection tasks.
+Moreover, we have provided a detailed analysis of four news headlines where ChatGPT correctly classified them using the prompt method but made errors when detecting clickbait.
+For example, given the English not-clickbait headline ”Tigers ’starved to death’ to make \$500 aphrodisiac wine with their bones,” ChatGPT misclassified it as clickbait.
+This is primarily because, first, the headline touches on sensitive topics such as animal abuse, illegal wildlife trade, and ethical controversies, which can easily elicit emotional responses from readers.
+Secondly, terms like ”starved to death” and ”aphrodisiac” are strong verbs and adjectives often used exaggeratedly or emotionally, and these words may match common vocabulary found in clickbait articles.
+Thirdly, in the absence of additional context, ChatGPT may be more inclined to categorize sentences with controversial or extreme characteristics as clickbait.For the Chinese not-clickbait headline ”扬州通报“网传副市长与副局长生活作风问题”:建议对二人免职” (Yangzhou reports ’alleged misconduct in the personal conduct of deputy mayor and deputy director’: Suggests dismissal of the two individuals.), ChatGPT misclassified it as clickbait.
+This is mainly because it is an attention-grabbing topic with relatively complex grammatical structures, including multiple parallel phrases and verb phrases.
+This complexity can lead the model to struggle in interpreting the sentence.
+It’s also important to note that the phrase ”网传” (spread on the internet) may have contributed to ChatGPT’s classification as spreading false information.For the English clickbait headline ”This is America’s favorite fast food restaurant,” ChatGPT misclassified it as not-clickbait.
+This is primarily because it is a topic related to American food culture and consumption habits, falling under the category of everyday life.
+Furthermore, the headline uses proper grammatical structure to pose a clear question, making it appear as a reasonable query rather than clickbait.For the Chinese clickbait headline ”与你有关!一批重要国家标准,今天发布!” (Related to You!
+A batch of important national standards released today!), ChatGPT misclassified it as not-clickbait.
+This is mainly because national standards and regulations are a topic of practical significance, and the proper grammatical structure is used to present a statement.
+These factors may have led ChatGPT to mistakenly classify it as a legitimate news headline rather than clickbait.
+In general, ChatGPT performs well in handling objective questions with clear, standardized answers.
+However, when it comes to addressing personalized questions, which often lack a fixed standard answer, it can be influenced by certain specific features.
+Moreover, conducting multiple tests on the same news may yield different results, leading to suboptimal performance in metrics like accuracy for clickbait detection.
+4.5 Ablation Study
+We compared the results between different prompts (as shown in Table 12) for clickbait detection tasks.
+As same as the main experiments, we select GPT-3.5 as a backbone for the prompts experiments.
+The results comparison of different Prompts are shown in Table 5.
+Generally speaking, there is no significant performance gap between the first prompt (P1) and the second prompt (P2) in terms of effectiveness for clickbait detection.
+Specifically, the performance of P1 is significantly better than that of P2 in the dataset WeChat.
+In dataset Sina, the opposite is true.
+Table 5: Comparison of different prompts for GPT-3.5(%). The bolder ones mean better.
 Data Method Shot
 5 10 20
+Comparison of different prompts for GPT3.5 on English datasets
 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
-CND GPT3.5+P1 80.75 69.06 78.69 73.56 79.33 69.51 70.08 69.79 82.07 75.66 70.08 72.76
+DLC GPT3.5+P1 80.75 69.06 78.69 73.56 79.33 69.51 70.08 69.79 82.07 75.66 70.08 72.76
 GPT3.5+P2 67.98 52.20 82.79 64.03 75.25 59.33 88.02 70.88 79.80 66.45 82.64 73.67
-SCC GPT3.5+P1 58.41 59.02 49.36 53.76 59.44 60.76 50.08 54.91 60.90 63.33 49.20 55.38
+CND GPT3.5+P1 58.41 59.02 49.36 53.76 59.44 60.76 50.08 54.91 60.90 63.33 49.20 55.38
 GPT3.5+P2 60.00 56.83 68.36 62.06 59.00 59.28 61.00 60.13 60.63 58.78 66.88 62.57
-DLC GPT3.5+P1 83.94 88.63 78.19 83.08 86.14 92.41 78.85 85.09 88.73 95.54 81.26 87.82
+SCC GPT3.5+P1 83.94 88.63 78.19 83.08 86.14 92.41 78.85 85.09 88.73 95.54 81.26 87.82
 GPT3.5+P2 81.49 76.54 90.01 82.73 84.39 82.71 86.93 84.77 87.52 87.30 87.81 87.55
-Table 8: Comparison of different prompts for GPT3.5 on Chinses datasets (%).
-Data Method Shot
-5 10 20
+Comparison of different prompts for GPT-3.5 on Chinese datasets
 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
@@
 GPT3.5+P2 68.13 72.18 59.59 65.28 66.43 69.54 55.77 61.9 70.94 72.14 67.21 69.59
-Evaluation metrics
-To test the effect of detection, we conduct four evaluation metrics to evaluate our method in experiments, such as accuracy, precision, recall and F1-score.
-accuracy The accuracy can be defined as the ratio of correctly predicted samples to the total number of samples.
-$Acc=\frac{TP+TN}{TP+TN+FP+FN}$
-precision The positive prediction rate can be defined as the ratio of correctly predicted positive samples to the total number of positive samples in the prediction.
-$P=\frac{TP}{TP+FP}$
-recall The positive precision can be defined as the ratio of correctly predicted positive samples to the total number of samples labeled positive.
-$R=\frac{TP}{TP+FN}$
-F1-score The F1-score can be defined as the harmonic average of precision and recall.
-$F1=\frac{{2}\times{P}\times{R}}{P+R}$
-4.4 Experimental Results
-The main results of the experiments on clickbait detection in English and Chinese datasets on zero-shot and few-shot scenarios are listed in Table 4, 5 and 6 respectively.
-We can see that the performance of ChatGPT has not demonstrated the best results on all four metrics over the seven datasets.
-Compared to the deep neural networks and fine-tuning PLMs, ChatGPT has a lot of room for improvement in clickbait detection.
-Moreover, with the increasing of pre-trained information, the effectiveness of clickbait detection, sometimes, is worse.
-Therefore, we can observe that a small amount of pre-training information does not significantly impact the performance of ChatGPT in detecting clickbait.
-Moreover, we can observe that ChatGPT can achieve stable results on almost all evaluation metrics.
-The results confirmed that LLMs can be adapted to other languages, and ChatGPT is a monolithic model capable of supporting multiple languages, which makes it a comprehensive multilingual clickbait detection technique.
-4.5 Ablation Study
-We compared the results between different prompts (as shown in Table 1 and Table 2) for clickbait detection tasks.
-As same as the main experiments, we select GPT3.5 as a backbone for the prompts experiments.
-The results comparison of different Prompts is shown in Table 7 and Table 8.
-Generally speaking, there is no significant performance gap between the first prompt (P1) and the second prompt (P2) in terms of effectiveness for few-shot clickbait detection.
-Specifically, the performance of P1 is significantly better than that of P2 in the dataset Wechat.
-In dataset Sina, the opposite is true.
+4.6 The Impact of Different LLMs
+Due to the rapid development and diverse characteristics of LLMs, we selected several different LLMs as the foundation for our experiments on clickbait detection.
+Specifically, we chose local LLMs including LLaMA3-8B, Yi-34B, and DeepSeek-32B, and compared their performance with commercial models such as GPT-3.5 and GPT-4.
+The results are presented in Tables 6 and 7.
+We observe that GPT-4 consistently performs well across most datasets.
+Among the local LLMs, DeepSeek-32B demonstrates relatively stable performance and even surpasses GPT-3.5 in certain metrics.
+However, it is important to note that some LLMs tend to favor specific classes, resulting in a significant imbalance between precision and recall.
+As a result, LLMs still show considerable limitations in the task of clickbait detection.
+Table 6: The Accuracy, Precision, Recall, F1-scores results (%) on English datasets with different LLMs(%). The bolder ones mean better.
+DLC CND SCC
+Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
+LLaMA3 77.68 66.41 69.67 68.00 59.32 61.52 49.77 55.02 82.95 88.78 75.43 81.56
+Yi 76.71 82.91 39.75 53.74 58.19 58.81 54.66 56.66 79.10 84.12 70.58 76.92
+DeepSeek 72.11 57.80 66.80 61.98 62.85 62.53 64.13 63.32 84.13 90.22 76.57 82.83
+GPT-3.5 80.75 69.06 82.79 73.56 60.00 59.02 68.36 62.06 83.94 88.63 90.01 83.08
+GPT-4 90.54 94.00 77.05 84.68 68.61 69.47 66.97 68.20 94.62 98.83 90.30 94.37
+Table 7: The Accuracy, Precision, Recall, F1-scores results (%) on Chinese datasets with different LLMs(%). The bolder ones mean better.
+Sina Tencent WeChat Paper
+Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1 Acc Pre Rec F1
+LLaMA3 73.52 69.12 85.02 76.25 63.87 60.04 82.98 69.67 54.64 52.91 84.52 65.08 70.63 64.42 92.16 75.83
+Yi 68.89 71.07 69.97 70.55 72.54 68.87 73.46 71.08 63.68 61.02 75.71 67.52 76.45 76.54 76.30 76.30
+DeepSeek 82.82 93.67 76.83 84.42 78.45 77.35 80.47 79.03 74.66 71.01 83.33 76.66 79.39 88.32 74.70 80.89
+GPT-3.5 70.73 76.21 60.28 67.32 59.34 57.91 71.13 63.84 60.12 56.69 85.71 68.24 68.33 72.18 62.70 66.42
+GPT-4 73.87 93.08 75.57 76.96 65.97 67.76 75.94 71.58 73.79 69.18 88.57 77.68 77.75 80.04 73.85 76.82
 5 Conclusion and Future Work
-In this paper, we present a study of the performance of LLMs (ChatGPT with GPT3.5) for clickbait detection.
+In this paper, we present a study of the performance of LLMs (ChatGPT with GPT-3.5 and GPT-4) for clickbait detection.
 During the benchmark experiments on both English and Chinese datasets, LLMs did not perform as well as the current state-of-the-art deep neural networks and fine-tuning PLMs methods in the realm of multilingual clickbait detection.